Friday, February 20, 2015

Night Log 2015-02-19

SA: Alexei
SO: Thea


=> Clear night with moderate seeing at the beginning and very bad seeing at the end. Were closed by humidity.

=> Data were taken for the following programs:
2014-2-SCI-017 P2 RSS (Optical preparatory survey for Cold HI 21-cm Absorption Line Survey (CHITALS))
2014-2-SCI-060 P3 RSS (Searching for binaries in unusual central stars of planetary nebulae)
2014-2-SCI-025 P3 RSS (Chemistry and kinematics of Blue Compact Dwarfs)
2014-1-MLT-007 P4 RSS (x2) (Long term monitoring of circumstellar disk in Be X-ray binaries)

=> Some calibration data for HRS

No comments:

Post a Comment